40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 29 21:43:27 2010, Alberto, Configuration, PSL, Reference Cavity PD Noise Spectrum 2010-03-29_FSS_PD_shotnoise_and_darknoise.png2010-03-29_FSS_PD_shotnoise_and_darknoise.png
    Reply  Tue Mar 30 06:37:32 2010, rana, Configuration, PSL, Reference Cavity PD Noise Spectrum rc.png
       Reply  Wed Mar 31 15:31:53 2010, steve, Update, PSL, Reference Cavity RF PD base upgraded brspdbs.JPG
    Reply  Sat Apr 3 11:35:47 2010, rana, Configuration, PSL, Reference Cavity PD Noise Spectrum 
       Reply  Sat Apr 3 16:07:40 2010, Alberto, Configuration, PSL, Reference Cavity PD Noise Spectrum 2010-03-29_FSS_PD_shotnoise_and_darknoise.png
          Reply  Tue Apr 6 11:27:11 2010, Alberto, Update, Computer Scripts / Programs, Data formats in the Agilent AG4395a Spectrum Analyzer 
             Reply  Tue Apr 6 16:55:28 2010, Alberto, Update, Computer Scripts / Programs, Data formats in the Agilent AG4395a Spectrum Analyzer 
Message ID: 2732     Entry time: Mon Mar 29 21:43:27 2010     Reply to this: 2733   2759
Author: Alberto 
Type: Configuration 
Category: PSL 
Subject: Reference Cavity PD Noise Spectrum 

[Rana, Alberto]

This evening we measured the noise spectrum of the reference cavity PD used in the FSS loop. From that we estimated the transimpedance and found that the PD is shot-noise limited. We also found a big AM oscillation in correspondence of the FSS modulation sideband which we later attenuated at least in part.

This plot shows the spectrum noise from the RF output of the photodetector.
 
 (here you should be able to see an attached figure, if not it's probably becasue imagemagic has having problems with displaying png files)
2010-03-29_FSS_PD_shotnoise_and_darknoise.png
 
The tall peak at 21.5 MHz is the AM modulation introduced by the EOM. It seems to be caused by a misalignment of the EOM which might be somehow modulating the polarization.
The mount in which the EOM sits is not very solid. We should change it with something similar to that of the other two EOMs in the Mach Zehnder.
By tightening down the plastic screws of the mount Rana reduced the amplitude of the AM modulation by 20dB.
 
The bump in both the dark and shot noise are in corrispondence of the resonance of the PD's electronics. As it appears, the electronics is not well tuned: the bump should coincide with the AM peak.
 
In the case of the dark noise spectrum, the bump is due to the thermal noise of the electronics. It's a good sign: it means that the electronics is good enough to be sensitive to it.
 
Transimpedance Estimate
As a "sanity check" we made an approximate estimate of the transimpedance to make sure that the PD is dominated by shot noise rather than other noises, ie electronic's noise.
 
  1. Supposing that the laser beam hitting the PD was shot noise limited, we measured 1.1V at the DC output. That let us estimate the photocurrent at DC of 20mA, for a 50Ohm output impedance.
  2. The shot noise for 20mA is 80 pA/rtHz
  3. From the nosie spectrum, we measured 3e-7 v/rtHz at 21.5MHz
  4. The impedance at RF is then Z_rf = 3e-7 V/rtHz / 80e-12 pA ~ 4000 Ohm
  5. Since the RF path inside the PD has a gain of 10, the transimpedance is ~400Ohm, which is about as we (ie Rana) remembered it to be.
  6. The PD seems to be working fine.
Attachment 2: 2010-03-29_FSS_PD_shotnoise_and_darknoise.png  40 kB  Uploaded Mon Mar 29 22:46:55 2010  | Show | Show all
ELOG V3.1.3-